### v1

# Explainable Machine Learning for ICU Readmission Prediction

Abstract
The intensive care unit (ICU) comprises a complex hospital environment, where decisions made by clinicians have a high level of risk for the patients’ lives. A comprehensive care pathway must then be followed to reduce patient complications. Uncertain, competing and unplanned aspects within this environment increase the overall difficulty in uniformly implementing the care pathway. Readmission contributes to this pathway’s difficulty, occurring when patients are admitted again to the ICU in a short timeframe, resulting in high mortality rates and high resource utilisation. Several works have tried to predict readmission through patients’ medical information. Although they have some level of success while predicting readmission, those works do not properly assess, characterise and understand readmission prediction. This work proposes a standardised and explainable machine learning pipeline to model patient readmission on a multicentric database (i.e., the eICU cohort with 166,355 patients, 200,859 admissions and 6,021 readmissions) while validating it on monocentric (i.e., the MIMIC IV cohort with 382,278 patients, 523,740 admissions and 5,984 readmissions) and multicentric settings. Our machine learning pipeline achieved predictive performance in terms of the area of the receiver operating characteristic curve (AUC) up to 0.7 with a Random Forest classification model, yielding an overall good calibration and consistency on validation sets. From explanations provided by the constructed models, we could also derive a set of insightful conclusions, primarily on variables related to vital signs and blood tests (e.g., albumin, blood urea nitrogen and hemoglobin levels), demographics (e.g., age, and admission height and weight), and ICU-associated variables (e.g., unit type). These insights provide an invaluable source of information during clinicians’ decision-making while discharging ICU patients.
keywords:
Readmission, Intensive Care Unit, Machine Learning, Explainable Predictions.
[…]
[label3]organization=Computational Biology and Clinical Informatics, Baker Heart and Diabetes Institute, city=Melbourne, state=Victoria, country=Australia
[…]
[label4]organization=Baker Department of Cardiometabolic Health, The University of Melbourne, city=Parkville, state=Victoria, country=Australia
[…]
[label7]organization=Faculty of Medicine, Dentistry and Health Sciences, The University of Melbourne, city=Parkville, state=Victoria, country=Australia
[…]
[label8]organization=The Department of Surgery, St. Vincent’s Hospital Melbourne, city=Fitzroy, state=Victoria, country=Australia
[…]
[label5]organization=Centre for the Digital Transformation of Health, School of Computing and Information Systems, The University of Melbourne, city=Parkville, state=Victoria, country=Australia
[…]
[label9]organization=The Department of Anaesthesia and Perioperative Medicine, Monash Health, city=Clayton, state=Victoria, country=Australia
[…]
[label10]organization=The Intensive Care Unit at Peninsula Private Hospital, Ramsay Healthcare, city=Langwarrin, state=Victoria, country=Australia
[…]
[label11]organization= The Department of Intensive Care Medicine, The Alfred Hospital, city= Prahran, state=Victoria, country=Australia
[…]
[label11b]organization= The Department of Anaesthesiology and Perioperative Medicine, The Alfred Hospital, city=Melbourne, state=Victoria, country=Australia
[…]
[label12]organization=The Australian and New Zealand Intensive Care Society Centre for Outcome and Resource Evaluation, city=Camberwell, state=Victoria, country=Australia
[…]
[label13]organization=The Australian and New Zealand Intensive Care Research Centre, School of Public Health and Preventive Medicine, Monash University, city=Melbourne, state=Victoria, country=Australia
[…]
[label14]organization= The Department of Intensive Care, Ballarat Base Hospita, city=Ballarat, state=Victoria, country=Australia
[…]
[label15]organization=The Department of Intensive Care, Austin Hospital, city= Heidelberg, state=Victoria, country=Australia
1 Introduction
Intensive care is a service for patients with likely recoverable and treatable conditions, providing them with more precise observation and invasive treatments Smith and Nielsen (1999). Appropriate patient management within an intensive care unit (ICU) is crucial. In a given ICU, the likelihood of long-term length of stay (LOS), organ failures, and mortality tends to increase if adequate management is not taken into consideration Moitra et al. (2016); Sakr et al. (2012); Ay et al. (2020). Clinicians who are caring for these patients also have multiple competing activities to consider Sujan et al. (2015); Turnbull et al. (2018), including the potential deterioration of the patient’s condition after discharge, emergency admissions, elective admissions for high-risk surgery, staffing and resources, and evidence for appropriate allocation of resources. These aspects highlight the challenges across an ICU setting.
Readmission is one of the factors that extends the challenges within an ICU setting. It occurs when patients are admitted again to the ICU in a short timeframe (between 48 hours and 30 days), resulting in high mortality rates, increased LOS and, consequently, high resource utilisation Mcneill et al. (2020); Alban et al. (2006); Kramer et al. (2013). In summary, when readmission occurs, it disrupts the care pathway for the patients and poses additional challenges to the clinical team and hospital caring for them.
Reducing ICU readmission rates might not only improve the care pathway, leading to better patient outcomes but also affect the hospital’s bottom line James (2013); Khera and Krumholz (2018). Predicting patients at high risk of readmission would not only allow early intervention to reduce the risk but also reduce mortality rates, reduce resource utilisation (based on LOS) and, potentially, hospital costs.
The emergence of abundant hospital electronic health record (EHR) data Yadav et al. (2018); Kim et al. (2019); Shickel et al. (2017) allowed the use of machine learning (ML) models targeting ICU readmission Badawi and Breslow (2012); Rojas et al. (2018); Mohanty et al. (2022); Thoral et al. (2021); McWilliams et al. (2019); Hammer et al. (2021); Lin et al. (2019); Xue et al. (2019); Barbieri et al. (2020); Hegselmann et al. (2022); Ashfaq et al. (2019). These readmission models have been summarised in Table S1. Apart from their undeniable importance, these models lack proper interpretation, generalisation and validation in different ICU settings (e.g., monocentric versus multicentric environments).
Our main contention is that predicting ICU readmission in a generalisable way with hospital-based datasets is still a challenge to overcome, even after taking into account these univariate predictive models and best practice ML techniques Lin et al. (2019); Hosein et al. (2013); Curto et al. (2016). The main reason for this difficulty is that medical data is usually noisy, uncertain, and characterised by a high degree of missingness, as it relies on different human inputs. Further exacerbating this complexity is data imbalance, with only a small proportion of patients being readmitted. ICU patients who happen to be readmitted are also very heterogeneous Hosein et al. (2013), meaning that finding novel patterns across an ICU readmission data cohort tends to be challenging. As a result, an appropriate translation from ML models to the clinical environment is still limited.
In addition, readmission’s association with healthcare quality has been debated in recent years Brown et al. (2014); Rewa et al. (2018). Many readmissions also emerge from factors inherent to patients, not only the quality of care, and may be reduced but not entirely eliminated. In these cases, better healthcare quality might not prevent readmission Pakbin et al. (2018). Still, predicting readmission accurately is valuable for post-discharge monitoring and planning. Machine learning models can connect clinical data to readmission, aiding pre-discharge decisions and informing readmission reduction programs Rojas et al. (2018); Thoral et al. (2021); Huang et al. (2021); Hu et al. (2022); McIlvennan et al. (2015).
This work proposes a standardised method for a comprehensive assessment of readmission modelling and prediction on ICU data Pollard et al. (2018, 2019) with explainable machine learning Molnar (2023). Our method resulted in a generalisable predictive model for both multicentric and monocentric data settings Pollard et al. (2018, 2019); Johnson et al. (2021). This is the first study to provide such a large and robust validation on readmission models. In addition, explanations from the constructed model also provided a better understanding of the main relevant patient variables associated with 30-day ICU readmission, which can potentially help in guiding clinicians while treating and discharging patients from the ICU.
1.1 State of Significance
Problem Intensive Care Unit (ICU) Readmission Prediction with Explainable Machine Learning.
What is Already Known Although there has been a range of related works aiming to predict ICU readmission with a variable level of success, those studies do not properly assess, characterise and understand readmission prediction, limiting their overall applicability to patient data.
What This Paper Adds This study proposes a standardised and explainable machine learning (ML) pipeline to model, predict and understand 30-day readmission. With a robust validation procedure, the developed ML model achieved generalisable performance on completely new patient data while deriving insightful (literature-supported) conclusions to contrast readmission and non-readmission within an ICU.
2 Material and Methods
Figure 1 presents the main methodological workflow followed in this work. Collecting data from monocentric (i.e., MIMIC IV) and multicentric databases (i.e., eICU) presents itself as the first step. Next, from both databases, the creation of variables describing the patient’s information takes place considering their admission timeframe. These variables include patient demographics, pre-admission information, vital signs and blood test results, and take into account the first and the last 24 hours, as well as their average values. Data preprocessing includes several steps, such as imputation, imbalance learning, and feature standardisation. The inclusion of variables outlining the hospitals and ICUs also plays an important role in this characterisation. Machine learning algorithms take these features as input to build a readmission model based on a timeframe of 30 days. The employed machine learning algorithms use multicentric data from eICU to build and internally validate the predictive model for readmission while it externally tests and validates the final model on monocentric data from MIMIC-IV. The readmission model predicts whether a given patient will be readmitted or not and supports explanations for its predictions, driving insights to clinicians while discharging a given patient in the ICU.
Figure 1: The methodological workflow followed by this work. Data comes from monocentric and multicentric databases (i.e., MIMIC IV and eICU, respectively), where the characterisation of ICU patients, hospitals, and ICU takes place. Preprocessing filters out, standardises and imputes data for the development of a 30-day readmission machine learning model, which is built and validated on eICU, and tested on MIMIC-IV. This model drives explanations from variables and predictions, potentially assisting and guiding clinicians while treating and discharging new patients from the ICU.
2.1 Data Sources
For assessing, analysing and understanding ICU readmission, we utilise two datasets: (i) the eICU Collaborative Research Database (eICU) Pollard et al. (2018, 2019); Goldberger et al. (2000), and (ii) the Medical Information Mart for Intensive Care (MIMIC IV) Johnson et al. (2021); Goldberger et al. (2000), which are described in next subsections. Their access relied on PhysioNet Goldberger et al. (2000).
2.1.1 eICU
The eICU contains data collected from 335 ICUs at 208 US hospitals in 2014 and 2015. This cohort has information about 166,355 unique patients and 200,859 admissions, accounting also for readmissions for those listed patients.
The eICU cohort is a de-identified multicentric dataset with information about patient demographics, admission and discharge details, APACHE IV scores and variables used for its calculation, vital signs, laboratory test results, medications, and care plan information.
2.1.2 MIMIC IV
The MIMIC cohort (version IV) emerged from two in-hospital systems of the Beth Israel Deaconess Medical Center (BIDMC): Electronic Health Record (EHR) and ICU-specific Clinical Information System. MIMIC IV has data about patients’ admissions through the emergency department or one of the intensive care units at BIDMC between 2008 and 2019. It includes data about 382,278 patients and 523,740 admission records.
Like eICU, this database is also de-identified and comprises demographics, hospital billing information, data related to patients’ admissions, transfers, and discharges, information about medication administration, and laboratory measurements. MIMIC IV also contains data generated in the ICU, such as intravenous and fluid inputs, procedures and services, charted information, and patients’ outputs.
2.2 Readmission Outcome
Patients were considered to have met the primary readmission outcome if they had one or more admission episodes within 30 days of discharge from their first ICU admission (see Supplementary Material and Methods for more details). If a patient is readmitted twice or more times, only the first readmission within 30 days counts for their data. It is worth noting that this range of days to define readmission follows different methodologies Mohanty et al. (2022); Lin et al. (2019); Xue et al. (2019); Barbieri et al. (2020).
Within the 30 days, the eICU cohort had a total of 149,009 patients, where 6,021 and 142,988 were admitted and non-readmitted, respectively. On the other hand, MIMIC IV had a total of 68,505 patients, where 5,984 underwent readmission and 62,521 did not.
2.3 Predictor Variable Selection
The same potential predictor variables were extracted from both datasets, which resulted in 168 common variables. Patient-level data included demographics (age, sex, weight, and height), admission diagnosis, highest and lowest physiological and biochemical values from the first and last 24 hours of ICU admission, treatments provided in ICU (ventilation, dialysis, and vasopressor) and overall illness severity scores (i.e., APACHE IV scores Lee et al. (2015)). Institutional-level variables included hospital-related (e.g., hospital size and location) and ICU-related variables (e.g., unit type and admission source). Supplementary Materials (Table S2) detail and distinguish these common variables in both databases across the following six categories: (i) demographics, (ii) vital sign- and blood test-related, (iii) pathology- and treatment-based, (iv) APACHE severity-related, (v) hospital-related, and (vi) ICU-related.
In addition, Figure S1 and Supplementary Material and Methods cover more details regarding the data collection criteria and the prediction window. Furthermore, categorical variables were encoded in a one-hot encoding scheme (a.k.a, one-of-K or dummy). Given the transformation in the categorical feature space, the number of features increased from 168 to 186 variables.
It is also worth noting that eICU contains an extra level of information on variables when compared to MIMIC IV. The variables from the aforementioned categories, which are detailed in Table S2, are only those that match across both cohorts or do not have linkage to the target readmission variable (e.g., length of stay or mortality). Nevertheless, as we used eICU to train and validate the machine learning models, we have included important eICU variables (missing in MIMIC IV) to analyse readmission. This step has been done only for the sake of readmission explainability. This list of complementary variables is detailed in Table S3.
2.4 Data Preprocessing
We performed a standardised approach to preprocess both datasets (i.e., eICU and MIMIC IV). This approach included basic steps to filter patients’ information and yield valid and reliable information for readmission prediction.
2.4.1 General Filtering
First, this step removed patients that are younger than 18 years old from the cohort. In addition, the proposed filter did not consider patients who died during their first ICU. Furthermore, it separated patients who spent less than four hours in the ICU from the datasets. This preprocessing step also excluded patients with abnormal values in the generated variables, such as having an admission weight greater than 260 kg or having an admission height greater than 240 cm. Supplementary Material and Methods and Figure S2 detail these patients’ inclusion and exclusion criteria on both eICU and MIMIC IV cohorts.
2.4.2 Handling Missing Data
The employed imputation method took into consideration a threshold of 20% of missingness, meaning that our machine learning pipeline will only use variables with less than 20% of missing values. The imputation method handled missing values by replacing them with population means (in the case of continuous variables) or a constant (in the case of categorical variables).
2.4.3 Dealing with Imbalance
Given the low proportion of readmitted patients in the ICU, the proposed pipeline considered several methods from the Imbalanced-learn toolbox Lemaître et al. (2017) to tackle the hardness of learning on imbalanced data Fernández et al. (2018). Nevertheless, random undersampling over the majority class (i.e., the non-readmitted class) yielded similar results to other methods, being then selected by its simplicity and performance.
Random undersampling took into accounting for the different data distributions, from 1:1 (1 readmitted patient to 1 non-readmitted patient) to 1:9 (one readmitted patient to 9 non-readmitted patients) to the real distribution (approximately one readmitted patient to 25 non-readmitted patients in eICU, and one readmitted patient to 10 non-readmitted patients in MIMIC IV).
2.4.4 Standardisation
In keeping with best practices to improve the predictive performance of variables Raschka et al. (2022), standardisation was undertaken by transforming each value of a given variable. On each variable value, standardisation subtracts it from the variable’s mean and then divides it by the variable’s standard deviation. Although standardisation is not a matter for some machine learning models, it is for several others. We therefore applied the same pipeline for all of them, making sure the use of proper data Raschka et al. (2022).
2.5 Machine Learning Modelling, Evaluation and Feature Selection
Machine learning (ML) modelling followed traditional development and evaluation guidelines Raschka (2018). The proposed modelling included the use of seven ensemble-based ML algorithms for patient readmission prediction: (i) explainable boosting machine (EBM) Nori et al. (2019), (ii) extremely randomised trees (ERT) Geurts et al. (2006), (iii) random forest (RF) Breiman (2001), (iv) extreme gradient boosting (XGBOOST) Chen and Guestrin (2016), (v) adaptive boosting (ADABOOST) Freund and Schapire (1995), (vi) gradient boosting (GB) Friedman (2001), and (vii) light gradient boosting machine (LightGBM) Machado et al. (2019). Table S4 presents the hyper-parameter optimisations for these ML algorithms.
It is important to emphasise that these algorithms have been chosen not only because of their robustness based on ensemble properties (i.e., by combining the results of several ML models) but also for their interpretation easiness. They all have (at some level) pre-built functions to describe feature importance for a given assessed dataset and can be straightforwardly used with other methods/tools for explainability.
With these algorithms as base classifiers, a bottom-up greedy feature selection ensued to ensure low ML complexity while maintaining generalisable predictive performance. Greedy feature selection starts its iterative process by having a bucket with zero features, adding one feature at each round. This method adds a feature in its bucket if it achieves the best predictive performance on a 10-fold cross-validation procedure, which relies on a model built by a classification algorithm. The evaluation of each feature relies on an average between Matthew’s correlation coefficient (MCC) Chicco and Jurman (2020) and the area under the receiver operating characteristic (AUC) curve Raschka et al. (2022); Raschka (2018); Sokolova and Lapalme (2009), which therefore considers resilience to class imbalance. In addition, other evaluation measures complemented the decision process in our ML pipeline, including the area under the precision-recall curve (APR), balanced accuracy, precision, recall (sensitivity), specificity, and F1 score Raschka et al. (2022); Raschka (2018); Sokolova and Lapalme (2009).
2.6 Statistical Analyses
Data are presented as number/proportion (percentage), mean (± standard deviation) or median (inter-quartile range) depending on the type and distribution of data. Chi-squared, ANOVA, Kruskal-Wallis and Wilcoxon rank-sum tests were used to compare groups depending on the type and distribution of data and the number of groups examined Demšar (2006). Statistical significance considers p-values lower than 0.001. The sample size of eICU and MIMIC IV datasets determined this decision, which consequently guaranteed the robustness of this study.
It is also worth noting that statistical analyses performed readmission comparisons not only within the same population (e.g., eICU) but also between populations (i.e., eICU versus MIMIC IV), strengthening this methodology. Furthermore, APACHE scores are provided only for eICU databases, while not directly available for MIMIC IV, enriching the explanation of characteristics of the former to the latter.
2.7 Model’s Diagnostic Interpretation, Analysis, and Explanations
The proposed ML pipeline for readmission also assesses its final model with a posteriori interpretation and explanation methods, which are described next.
2.7.1 Likelihood Ratios
Likelihood ratio (LR) McGee (2002) provides a measure for expressing the diagnostic accuracy of a clinical finding of the readmission model from this work. LR represents the probability of patients with a condition or undergoing a medical event (e.g., being readmitted) divided by the probability of patients without a condition or undergoing a medical event (e.g., not being readmitted). Positive LRs in readmission prediction estimate the model’s capabilities of identifying readmitted against non-readmitted patients. When LRs are greater than one, it means they go along with the readmission diagnosis. In other words, the greater the LR is, the more certain the predictive model is in finding readmitted patients against non-readmitted patients.
2.7.2 Calibration
The calibration technique visually evaluates the predictive readmission model by inspecting the calibration curves and also numerically estimating the calibration slope and intercept, as well as the integrated calibration index (ICI) and associated metrics (E50, E90, Emax ) Austin and Steyerberg (2019). The ideal calibration slope is equal to one. The ideal calibration intercept, as well as the ideal value for the ICI and its related metrics, is zero.
2.7.3 Shapley Additive Explanations
Our pipeline also utilised Shapley Additive Explanations (SHAP) Lundberg and Lee (2017), which is a game theory-based method to derive explanations for individual predictions Molnar (2023), in which combined results yield a global interpretation of the model’s behaviour. In readmission prediction, SHAP assists with model analysis, leading to a better understanding of readmission through a range of explanations. As a result, it can be used to recommend a set of personalised ICU guidelines while treating/discharging patients in the ICU.
3 Results
3.1 eICU Readmission Group Characteristics, Outcomes and Comparison to MIMIC IV
Table 1 provides the analyses contrasting readmitted and non-readmitted population groups. Out of 149,009 ICU admissions in the eICU database, readmission happens in 6,021 (4.04%) patients within 30 days of their first ICU admission. We have observed that the readmitted group includes older, sicker and more male patients. These results are corroborated by a statistical test with a significance level lower than 1%. The readmitted group also had a high level of gastrointestinal, respiratory and sepsis admission diagnoses when compared to the non-readmitted group (with a p-value <<< 1%).
The combined medical-surgical ICU and cardiothoracic surgical ICU types reached the lowest proportion of readmissions, with statistical significance. The readmitted group received more vasopressors and renal replacement therapies when contrasted with the non-readmitted group. The readmitted group commonly had a longer ICU stay and, consequently, a longer hospital length of stay (p-value <<< 1%). Finally, they were more than twice as likely to die in a hospital (16.0% versus 7.1% with a p-value <<< 1%), with survivors needing more rehabilitation, home nursing or skilled nursing facilities (p-value <<< 1%).
The Medical-Surgical Intensive Care Unit (MSICU) had the highest readmission rate at 49.1%, followed by the Medical Intensive Care Unit (MICU) with 10.9% and the Critical Care Unit-Cardio-Thoracic Intensive Care Unit (CCU-CTICU) with 10.1%. However, the absolute difference between readmission and non-readmission rates in most ICU types does not exceed 2.7%, except for MSICU, with a 6.6% difference between the two groups. The lowest readmission rate has been found in the Cardio-Surgical Intensive Care Unit (CSICU).
Table 1: Characteristics of discharged readmitted and non-readmitted patients in eICU.
Characteristic
Readmitted
(value with % OR
mean/median)
Non-Readmitted
(value with % OR
mean/median)
p-value
Readmission Numbers 6,021 142,988 -
Gender:
Male 3,417 (56.8%) 77,144 (54.0%) 2.198e-05*
Female 2,602 (43.2%) 65,771 (46.0%) 2.198e-05*
Age 65.48 / 67.00 62.70 / 65.00 <2.2e-16*
Body Mass Index (BMI) 28.69 / 27.30 29.05 / 27.51 <0.001*
APACHE IV 59.11 / 56.00 53.86 / 50.00 <2.2e-16*
APACHE Diagnosis:
Cardiovascular 1,789 (29.7%) 45,423 (31.8%) <0.001*
Neurological 881 (14.6%) 27,068 (18.9%) <2.2e-16*
Gastrointestinal 798 (13.3%) 13,755 (9.6%) <2.2e-16*
Trauma 230 (3.8%) 6,470 (4.5%) 0.011
Respiratory 1,045 (17.4%) 22,158 (15.5%) <0.001*
Sepsis 972 (16.1%) 17,404 (12.2%) <2.2e-16*
Other 412 (6.8%) 13,810 (9.7%) 3.8e-13*
Unit Type:
Critical Care Unit-Cardio-Thoracic Intensive Care Unit (CCU-CTICU) 610 (10.1%) 12,322 (8.6%) 4.8e-05*
Cardio-Surgical Intensive Care Unit (CSICU) 90 (1.5%) 5,335 (3.7%) <2.2e-16*
Cardio-Thoracic Intensive Care Unit (CTICU) 291 (4.8%) 4,750 (3.3%) 2.7e-10*
Cardiac Intensive Care Unit (CICU) 440 (7.3%) 9,392 (6.6%) 0.025
Medical Intensive Care Unit (MICU) 657 (10.9%) 11,710 (8.2%) 7.6e-14*
Surgical Intensive Care Unit (SICU) 503 (8.4%) 8,854 (6.2%) 1.5e-11*
Medical-Surgical Intensive Care Unit (MSICU) 2,955 (49.1%) 79,613 (55.7%) <2.2e-16*
Neurological Intensive Care Unit (NICU) 475 (7.89%) 11,012 (7.7%) 0.610
Hospital Capacity:
Extra-Large Hospital
(>500 beds)
2,427 (40.3%) 51,538 (36.0%) 1.7e-11*
Large Hospital
(250 – 499 beds)
1,431 (23.8%) 33,108 (23.2%) 0.277
Medium-sized Hospital
(100 – 249 beds)
1,052 (17.5%) 31,539 (22.1%) <2.2e-16*
Small hospital
(<100 beds)
231 (3.8%) 9,438 (6.6%) <2.2e-16*
Unknown number of beds 880 (14.6%) 17,365 (12.1%) 1.1e-08*
Hospital Type:
Teaching 1,687 (28.0%) 35,715 (25.0%) 1.1e-07*
Non-teaching 4,334 (72.0%) 107,273 (75.0%) 1.1e-07*
Patient Origin:
ICU 152 (2.5%) 3,261 (2.3%) 0.232
Operating Room 564 (9.4%) 14,675 (10.3%) 0.026
Emergency Department (ED) 2,089 (34.7%) 59,728 (41.8%) <2.2e-16*
Recovery Room 222 (3.7%) 4,929 (3.4%) 0.3359
Unknown 1,545 (25.7%) 33,334 (23.3%) 2.7e-05*
Other 1,449 (24.1%) 27,061 (18.9%) <2.2e-16*
Therapies in ICU:
Mechanical Ventilation 1,252 (20.8%) 27,651 (19.3%) 0.005
Vasopressors 761 (12.6%) 16,060 (11.2%) <0.001*
Renal Replacement Therapies (RRT) 232 (3.9%) 4,406 (03.1%) <0.001*
Length of Stay (LOS):
ICU LOS, days 3.72 / 2.18 2.86 / 1.75 <2.2e-16*
Hospital LOS, days 17.86 / 14.64 6.92 / 5.00 <2.2e-16*
Hospital Mortality 965 (16.0%) 10,162 (7.1%) <2.2e-16*
Discharge destination:
Home 2,290 (38.0%) 87,841 (61.4%) <2.2e-16*
Care Facility 2,409 (40.0%) 38,220 (26.7%) <2.2e-16*
Care Facility (Discharge destination):
Skilled Nursing Facility 1,150 (19.1%) 18,578 (13.0%) <2.2e-16*
Other Hospital 374 (6.2%) 5,698 (4.0%) <2.2e-16*
Rehabilitation 428 (7.1%) 6,213 (4.3%) <2.2e-16*
Nursing Home 128 (2.1%) 1,604 (1.1%) 1.7e-12*
Other External 329 (5.5%) 6,127 (4.3%) 1.2e-05*
$\star$ Presence of statistical significance.
Figure S3 also highlights patients readmitted to another type of ICU where the patient was originally admitted. Several ICUs in the eICU database have more than 30% readmissions to other ICU types, such as the Surgical Intensive Care Unit (SICU), Medical Intensive Care Unit (MICU) and Cardiac Intensive Care Unit (CICU).
While aiming to compare MIMIC IV and eICU readmitted patients, we noticed several differences in the level of information these cohorts provide, including patients who do not have length of stay (LOS) and discharge location. As a result, MIMIC IV and eICU were reduced from 5,984 and 6,021 to 5,980 and 5,403, respectively. This step was performed only to compare both cohorts directly. Matching the characteristics of MIMIC IV and eICU is not straightforward overall. MIMIC IV does not include several characteristics contained in eICU, and vice-versa, meaning that the comparison across the readmission groups from eICU and MIMIC IV was restricted to only a few characteristics.
Apart from that, we could observe they have similar baseline characteristics (Table 2). However, MIMIC IV readmitted patients received more ventilation and vasopressors but fewer renal replacement therapies than the readmitted eICU patients. In addition, MIMIC IV readmitted patients stayed longer in the ICU and hospital stay. Mortality was also slightly higher in the MIMIC IV readmission patients, although no statistical difference has been found between the two databases. Among survivors, the readmission group in the eICU was discharged more to home, while readmitted patients in MIMIC IV had a higher rate of further treatments in a rehabilitation care facility (p-value <<< 1%).
Table 2: Comparisons of characteristics between eICU and MIMIC IV readmission groups.
Characteristic
eICU
(value with % OR
mean/median)
MIMIC
(value with % OR
mean/median)
p-value
Readmission Numbers 5,403 5,980 -
Gender:
Male 3,077 (56.9%) 3,382 (56.6%) 0.700
Female 2,326 (43.1%) 2,598 (43.4%) 0.685
Body Mass Index (BMI) 28.7 / 27.3 28.9 / 27.8 0.010
Age 65.5 / 67.0 65.7 / 67.0 0.620
Patient Origin:
Emergency Department (ED) 1,882 (34.8%) 2,824 (47.2%) <2.2e-16*
Therapies in ICU:
Mechanical Ventilation 1,146 (21.2%) 1,410 (23.6%) 0.003
Vasopressors 703 (13.0%) 1,111 (18.6%) 6.6e-16*
Renal Replacement Therapies (RRT) 215 (4.0%) 66 (1.1%) <2.2e-16*
Length of Stay (LOS):
ICU LOS, days 3.7 / 2.2 3.5 / 2.0 6.2e-05*
Hospital LOS, days 17.9 / 14.6 20.2 / 15.0 9.5e-06*
Hospital Mortality 882 (16.3%) 1,089 (18.2%) 0.009
Discharge destination:
Home 2,055 (38.0%) 631 (10.6%) <2.2e-16*
Care Facility (Rehabilitation) 407 (7.5%) 697 (11.7%) 1.5e-13*
Care Facility (Skilled Nursing Facility) 1,052 (19.5%) 1,132 (18.9%) 0.479
$\star$ Presence of statistical significance.
3.2 Machine Learning Validation
The statistical analysis provided an insightful characterisation of the readmission group, contrasting it with the non-readmission group and across populations. With this information, we now move to a machine learning analysis and validation.
Table S5 shows the 47 selected features by the greedy feature selection approach, considering a balanced readmission dataset. The details of these variables representing machine learning features can be found in Table S2. Although we employed several different undersampling techniques to tackle the high level of readmission imbalance, using a balanced readmission rate presented itself as the most successful technique, which was then applied in both eICU and MIMIC IV cohorts at first. Hyper-parameter optimisation has also been applied after feature selection, aiming to improve the predictive performances of the resultant models (Table S4).
When bringing all these ML components together, our proposed pipeline achieved an area under the ROC curve (AUC) of 0.68 on multicentric eICU data under a 10-fold cross-validation procedure with a Random Forest classifier with 80 trees in a balanced readmission scenario. On the blind test (stratified 10% from eICU data), which internally validates the proposed model, consistent results were reached (AUC of 0.672) when compared to 10-fold cross-validation. When externally validated on MIMIC IV’s data, our model reached an AUC of 0.616, which demonstrates the overall generalisation capabilities of the proposed model to predict 30-day readmission. Figure 2 shows the AUC plots for these three validation schemes. The results for other performance measures are also summarised in Table S6.
Figure 2: The predictive performance of the proposed readmission model on 10-fold cross-validation and blind test on eICU data. External validation was made utilising MIMIC IV data.
3.3 Calibration Analysis
Calibration metrics are presented in Table S7 for our readmission model on 10-fold cross-validation and blind testing on eICU, and external validation on MIMIC IV. Given these metrics, we present the calibration curves for these evaluation sets in Figures S4-S6.
Figure S4 depicts the calibration curve for our proposed readmission model on 10-fold cross-validation on eICU data. The calibration curve on the results in cross-validation indicates a slight overestimation of risk at predicted probabilities between 0.2 and 0.4, and a slight underestimation of risk at predicted probabilities between 0.6 and 0.7. Figure S5, in turn, shows the calibration curve for our model on a blind test set from eICU data. The calibration curve indicates a slight overestimation of risk at predicted probabilities around 0.5. Finally, Figure S6 outlines the calibration curve for our model on an external validation set, i.e., MIMIC IV. Its calibration curve indicates a slight underestimation of risk at predicted probabilities below 0.2, and an overestimation of risk at predicted probabilities above 0.4 but most significantly above 0.6. Our main conclusion is that our model is well-calibrated in all sets, demonstrating its clinical usability in an ICU setting.
3.4 Likelihood Ratio Analysis
Figure S7 summarises the likelihood ratios (LRs) for our readmission model, which performs using a classification probability threshold of 0.5. With this threshold, the model’s LRs for cross-validation on eICU, blind testing on eICU, and external validation on MIMIC IV are 1.660 (estimated change of probability of 9.628%), 1.636 (estimated change of probability of 9.348%) and 1.335 (estimated change of probability of 5.494%), respectively. These model’s results indicate its reasonable diagnostic expression.
Based on Figure S7, a threshold between 0.70 and 0.82 benefits the model the most regarding the LRs. The local and global optima based on the classification threshold 10-fold cross-validation would be 0.712 and 0.813, respectively. With 0.712, the model’s LR for 10-fold cross-validation on eICU, blind testing on eICU and external validation on MIMIC IV would be 2.905 (estimated change of probability of 20.26%), 3.326 (estimated change of probability of 22.831%) and 1.942 (estimated change of probability of 12.615%), respectively. These estimated changes show small but relevant shifts in the post-test probabilities.
With 0.813, the model reaches a maximum LR on 10-fold cross-validation on eICU (i.e., 4.256), resulting in a change of probability of readmission for 27.517%. Nevertheless, with this threshold, we have a lower LR for blind testing and external validation. In the blind test on eICU, an LR of 1.075 would be identified (with an estimated change of probability of 1.373%), while for the external validation on MIMIC IV, the model would result in an LR of 1.400 (with an estimated change of probability of readmission of 6.393%).
This analysis determines that setting properly the classification threshold is beneficial for having a better diagnostic readmission profile for ICU patients. In summary, the model’s diagnostic power depends on a proper analysis comparing and contrasting machine learning and clinical predictiveness.
3.5 Explainable Machine Learning for ICU Readmission
We used SHAP to demonstrate how individual features included in the model (see Tables S2 and S5) influence overall readmission predictions. Figure 3 summarises and highlights the 20 most important features, based on SHAP values, we employed to characterise ICU readmission in eICU and MIMIC IV.
Figure 3: The SHAP summary plot for our proposed readmission model on eICU training data. We show the 20 features with higher SHAP values, i.e., that have a higher impact on the model’s predictive outputs.
Most of the features in the SHAP plot of Figure 3 are vital sign- and blood test-related (see Table S2). The three most important in this category are the minimum value of albumin in the blood during the first 24 hours, the maximum level of blood urea nitrogen (BUN) during the last 24 hours and the minimum level of hemoglobin during the last 24 hours.
We identified through the SHAP tree explanation model that high values for the variables Min. Albumin (First 24hs) and Min. Hemoglobin (Last 24hs) are usually more linked to the non-readmission of patients, while low values for them are more present in readmitted patients. Max. Blood Urea Nitrogen (Last 24hs) reveals an opposite trend, where higher values for this variable are fairly more associated with ICU patient readmission, and there is clear evidence that low values for the Max. Blood Urea Nitrogen (Last 24hs) result are attached to non-readmission.
In addition, the SHAP tree explanation model resulted in only one hospital-related variable (see Figure 3 and Table S2), i.e., the number of minutes from unit admission time that the patient has been admitted to the hospital (i.e., the Hospital Admission Offset). In this case, there is a high correlation between lower values of Hospital Admission Offset and non-readmission, whereas relatively higher values for this variable are more related to readmission.
From ICU-related variables, the SHAP model outputs a category of unit type as important – i.e., the Unit Type Med-SurgICU, which stands for a unit type of Medical-Surgical ICU. As the variable Unit Type corresponds to a categorical variable and Med-SurgICU is one of its options, we transformed its value via one-hot encoding. In the SHAP summary plot of Figure 3, this basically means that most patients coming from Med-SurgICU are not facing readmission, while patients that are not passing through such unit type undergo readmission.
Three important demographics-related variables were also tagged in the SHAP summary from Figure 3: Age, Admission Weight and Admission Height. As seen in this figure, older people were often associated with readmission, while low values in age are noticed in non-readmitted patients. This factor is actually expected as elderly patients tend to present more comorbidities Divo et al. (2014), resulting in more readmissions. Contrarily, admission weight followed a distinct pattern, where readmitted patients had low-medium values in their readmission weight. We also noticed that readmission and non-readmission were found to be more present in high and low values of the variable admission height, respectively.
Finally, pathology-related variables, such as the Min. Partial Thromboplastin Time (First 24hs), are also at the top 20 features ranked by SHAP values. As seen in Figure 3, high Min. Partial Thromboplastin Time (First 24hs) values are more connected with readmission.
We believe these explanations, derived by SHAP for the proposed readmission model, regarding the top 20 most impactful variables might serve as a new source of guidelines while assessing patients through ICU, especially while discharging patients. Although clinical practice and experience will always be compulsory in such conditions, we believe machine learning models might provide proper support for them.
4 Discussion
4.1 Key Findings
Although it is challenging to map and match common variables or characteristics across different cohorts, our study drove important clinical insights while analysing eICU and comparing it to MIMIC IV (Section 3.1). We discovered how heterogeneous the readmission monocentric and multicentric populations are in general. For example, the readmitted patients in MIMIC IV were treated more frequently with ventilation and vasopressors, although renal replacement therapies were more found in patients readmitted to eICU. Such contrasts in the cohort populations highlight the ability of our standardised pipeline to deal with heterogeneous readmission data.
Our proposed machine learning pipeline learned proper patterns on multicentric ICU data, consequently generalising well on an independent blind test set over the same data (Section 3.2). Our proposed machine learning model captured the essence of readmission, being able to transfer similar predictive performance to external validation on monocentric ICU data (Section 3.2). Overall, this shows learning on data coming from multiple ICUs has its limitations but may lead to generalisable predictive performance if done adequately.
A reasonable calibration level is also an essential aspect of the proposed readmission model (Section 3.3). As a result, our proposed model shares meaningful clinical decision-making due to its ability to calculate individualised probabilistic estimates of readmission Munn et al. (2022). For example, clinicians may decide that patients with a risk of readmission at least double that of the average risk in their ICU population should be flagged for routine monitoring following discharge from ICU. If the patient is moved to the ward, this may take the form of a brief clinical review by an intensivist each day following transfer to identify early signs of clinical deterioration. This could potentially facilitate the escalation of ward-based measures to optimise the patient’s management and avoid ICU readmission, or it could facilitate better preparation and planning such that unplanned readmissions are minimised.
A good diagnostic level, when we increase the classification probability threshold, complements the model’s predictive performance (Section 3.4). At the threshold(s) of 0.712 (and 0.813), our model returns the most reliable likelihood ratio among cross-validation, blind testing and external validation, yielding trustworthy readmission predictions.
Finally, we highlight how the 20 most impactful features are linked with readmission prediction through their explainability (Section 3.5). These explanations might be used to support the clinical team. In fact, some of the insights provided by the interpretation of the readmission model’s explanation are corroborated by the literature. We have highlighted these analyses and observations for some of the important variables displayed by SHAP values (Figure 3) next:
  • 1.
    Low serum albumin levels in ICU patients have already been linked with death, length of stay, and readmission Rojas et al. (2018); Herrmann et al. (1992). These results from previous work are congruent with our analysis based on Figure 3.
  • 2.
    High values of blood urea nitrogen (BUN) have also been associated with readmission and mortality in past studies Rojas et al. (2018); Arihan et al. (2018); Li et al. (2021), representing an important feature to characterise critically ill patients admitted to ICU.
  • 3.
    Previous studies also mapped low hemoglobin levels preceding ICU discharge, which may provide important insights for avoiding readmission Hammer et al. (2021); Warner et al. (2022); Kerfeld et al. (2020).
  • 4.
    Elderly patients usually are observed to be more associated with ICU readmission Lee et al. (2022); Glans et al. (2020); Pedersen et al. (2017).
  • 5.
    The patients’ weight is strongly correlated to readmission. While gaining weight through an ICU readmission is linked to non-readmission in patients with heart failure Howie-Esquivel et al. (2019), weight and weight change are common factors associated with mortality, duration of mechanical ventilation, and length of stay in the ICU You et al. (2013); Mishra et al. (2021). Therefore, looking at the patient’s admission weight and tracking their respective weight during ICU admission is a relevant aspect before discharging a patient from the ICU.
Overall, this shows our proposed model was able to learn and capture good insights on eICU data, being compatible with results and discussion in the ICU readmission literature.
4.2 Strengths and Limitations
One of the limitations of this work is the lack of comparable variables between multicentric eICU and monocentric MIMIC IV, including complete APACHE IV scores, which are not present in MIMIC IV. Nevertheless, we matched 168 common variables between the two cohorts, constituting both patient- and institutional-level information, building a key strength of our characterisation.
Another aspect is that several variables presented a large portion of missing data (more than 80%), bringing an extra level of difficulty to perform our methodology. We handled missing data by excluding columns with more than 20% of missing values and performing imputation in the other cases. This decision was taken because we considered we would include a high proportion of bias and noise if we impute columns with a great frequency of missingness. Although missingness might be informative Tan et al. (2023), the generalisation reached by the proposed predictive readmission model showed the robustness of our ensemble-based methodology to deal with different noise levels – from both missingness and utilised imputation approach. We plan in future work to better understand missingness across eICU and MIMIC IV as a way to improve machine learning modelling.
In addition, we observed the challenge of finding comparable metrics for the severity of the illness to describe the population between eICU and MIMIC. Therefore, our attempt resumed using the variables included in most severity of illness scores in our data description. Given data restrictions, we believe this was the best way to compare the severity of illness between both cohorts.
Finally, we noticed an overall heterogeneity of the studied populations, including imbalance due to rare occurrences of readmission, where the eICU cohort presents 4.04% of readmission cases while MIMIC IV has more than double (8.74%). We successfully dealt with the imbalance by employing appropriate performance metrics on top of the feature selection method Chicco and Jurman (2020). Our choice for the set of metrics considered the four confusion matrix categories (true positives, false negatives, true negatives, and false positives), setting proportional weights based on the size of the readmission group and the size of the non-readmission group in the datasets. Undersampling techniques also assisted with diminishing the imbalance hardness in our study.
4.3 Relationship to Previous Studies
Multiple predictive models were developed for readmission with different AUC performances (see Table S1). Nevertheless, we believe that most of these models lack clinical applicability. The main reason behind this is that those related works prefer highlighting their good predictive performances rather than understanding readmission well and bridging meaningful clinical and machine learning knowledge.
Although performing similarly to the literature, most of the works are focused on monocentric data, except for Badawi and Breslow Badawi and Breslow (2012) and Mohanty et al. Mohanty et al. (2022). However, both works lack clinical decision-making and are not well-validated with internal and external data sets. Apart from Mohanty et al. employing explainable machine learning in their proposed models, providing insights for their results, they remain local to the Epic EHR database employed by them.
Hegselmann et al.’s work Hegselmann et al. (2022) is the closest to our method. Their readmission model is built from University Hospital Münster (UKM) data and externally tested on MIMIC IV. Even though it presents a certain level of similarity to our work, their work is set in a controlled scenario, where UKM data is employed to build the model. Furthermore, both UKM and MIMIC IV are monocentric, presenting, therefore, a lower level of difficulty than multicentric eICU. In our method, we learn a readmission model from multicentric data to generalise on monocentric data. We show our method works well in translating from one to another.
Funding
D.B. A. was supported by an Investigator Grant from the National Health and Medical Research Council (NHMRC) of Australia (GNT1174405 to D.A.). Supported in part by the Victorian Government’s Operational Infrastructure Support Program.
Conflict of Interest
None declared.
Contributions
Alex G. C. de Sá: Conceptualisation, Draft figure proposals, Literature review, Methodology, Machine Learning - Experimental planning and execution, and results analysis, Visualisation, Writing - Original draft preparation and supplementary materials, Reviewing, Editing, and Validation (machine learning perspective);
Daniel Gould: Conceptualisation, Literature review, Visualisation, Investigation, Writing - introduction and results, Analysis, Reviewing, Editing, and Validation (clinical and machine learning perspectives);
Anna Fedyukova: Data collection, Machine Learning - Experimental planning and execution, and results analysis, Visualisation, Writing- Methodology and results, supplementary materials, and Validation (machine learning perspective);
Mitchell Nicholas: Literature review, Reviewing, Editing, and Validation (clinical perspective);
Calvin Fletcher: Writing - introduction, Validation (clinical perspective), Reviewing, and Editing;
Lucy Dockrell: Writing, Validation (clinical perspective), Reviewing, and Editing;
David Pilcher: Supervision, Writing, Validation (clinical perspective), Reviewing, and Editing;
Daniel Capurro: Conceptualisation, Supervision, Writing, Reviewing and Editing;
David B. Ascher: Supervision, Writing, Reviewing, Editing and Validation (clinical and machine learning perspectives);
Khaled El-Khawas: Conceptualisation, Supervision, Supervision, Writing- Methodology and Results, and Validation (clinical perspective);
Douglas E. V. Pires: Conceptualisation, Supervision, Writing, Reviewing, Editing, and Validation (clinical and machine learning perspectives);
Supplementary Information
Supplementary Materials are available at https://tinyurl.com/36f333sc.

### v4

@@
 Abstract
+Background
 The intensive care unit (ICU) comprises a complex hospital environment, where decisions made by clinicians have a high level of risk for the patients’ lives.
-A comprehensive care pathway must then be followed to reduce patient complications.
-Uncertain, competing and unplanned aspects within this environment increase the overall difficulty in uniformly implementing the care pathway.
-Readmission contributes to this pathway’s difficulty, occurring when patients are admitted again to the ICU in a short timeframe, resulting in high mortality rates and high resource utilisation.
+Uncertain, competing and unplanned aspects within the ICU environment increase the difficulty in uniformly implementing the care pathway.
+Readmission contributes to this pathway’s difficulty, resulting in high mortality rates and resource utilisation.
+Objectives
 Several works have tried to predict readmission through patients’ medical information.
 Although they have some level of success while predicting readmission, those works do not properly assess, characterise and understand readmission prediction.
-This work proposes a standardised and explainable machine learning pipeline to model patient readmission on a multicentric database (i.e., the eICU cohort with 166,355 patients, 200,859 admissions and 6,021 readmissions) while validating it on monocentric (i.e., the MIMIC IV cohort with 382,278 patients, 523,740 admissions and 5,984 readmissions) and multicentric settings.
-Our machine learning pipeline achieved predictive performance in terms of the area of the receiver operating characteristic curve (AUC) up to 0.7 with a Random Forest classification model, yielding an overall good calibration and consistency on validation sets.
-From explanations provided by the constructed models, we could also derive a set of insightful conclusions, primarily on variables related to vital signs and blood tests (e.g., albumin, blood urea nitrogen and hemoglobin levels), demographics (e.g., age, and admission height and weight), and ICU-associated variables (e.g., unit type).
-These insights provide an invaluable source of information during clinicians’ decision-making while discharging ICU patients.
+This work proposes a standardised and explainable machine learning pipeline to model, predict and explain patient ICU readmission.
+Method
+This work focuses on using machine learning to model readmission using a multicentric eICU database while validating it on both MIMIC-IV monocentric and eICU multicentric settings.
+SHAP (Shapley Additive Explanations), calibration and likelihood ratio analysis are employed to comprehend the behaviour of the predictive model.
+Results
+Our machine learning pipeline achieved predictive performance in the area of the receiver operating characteristic curve (AUC) up to 0.7 with a Random Forest classification model, yielding an overall consistency and generalisation on validation sets.
+After an interpretation analysis of our model through likelihood ratios and calibration, we observed that the proposed methodology could generate predictive models with proper and translatable diagnosis capabilities.
+In addition, from explanations provided by the constructed model, we could also derive a set of insightful conclusions, primarily on variables related to vital signs and blood tests (e.g., albumin, blood urea nitrogen and haemoglobin levels), demographics (e.g., age, and admission height and weight), and ICU-associated variables (e.g., unit type).
+Conclusions
+The model’s predictions, behaviour and explainability method provided insights that yielded valuable information, which clinicians might (be able to) use to make decisions while discharging ICU patients.
 keywords:
 Readmission , Intensive Care Unit , Machine Learning , Explainable Predictions.
+journal: International Journal of Medical Informatics
 […]
@@
 […]
-[label14]organization= The Department of Intensive Care, Ballarat Base Hospita, city=Ballarat, state=Victoria, country=Australia
+[label14]organization=The Department of Intensive Care, Ballarat Base Hospital, city=Ballarat, state=Victoria, country=Australia
 […]
@@
 1 Introduction
-Intensive care is a service for patients with likely recoverable and treatable conditions, providing them with more precise observation and invasive treatments Smith and Nielsen (1999).
+Intensive care is a service for patients with likely recoverable and treatable conditions, providing them with more precise observation and invasive treatments [1].
 Appropriate patient management within an intensive care unit (ICU) is crucial.
-In a given ICU, the likelihood of long-term length of stay (LOS), organ failures, and mortality tends to increase if adequate management is not taken into consideration Moitra et al.
-(2016); Sakr et al.
-(2012); Ay et al.
-(2020).
-Clinicians who are caring for these patients also have multiple competing activities to consider Sujan et al.
-(2015); Turnbull et al.
-(2018), including the potential deterioration of the patient’s condition after discharge, emergency admissions, elective admissions for high-risk surgery, staffing and resources, and evidence for appropriate allocation of resources.
+In a given ICU, the likelihood of long-term length of stay (LOS), organ failures, and mortality tends to increase if adequate management is not taken into consideration [2, 3, 4, 5].
+Clinicians who are caring for these patients also have multiple competing activities to consider [6, 7], including the potential deterioration of the patient’s condition after discharge, emergency admissions, elective admissions for high-risk surgery, staffing and resources, and evidence for appropriate allocation of resources.
 These aspects highlight the challenges across an ICU setting.
 Readmission is one of the factors that extends the challenges within an ICU setting.
-It occurs when patients are admitted again to the ICU in a short timeframe (between 48 hours and 30 days), resulting in high mortality rates, increased LOS and, consequently, high resource utilisation Mcneill et al.
-(2020); Alban et al.
-(2006); Kramer et al.
-(2013).
+It occurs when patients are admitted again to the ICU in a short timeframe (between 48 hours and 30 days), resulting in high mortality rates, increased LOS and, consequently, high resource utilisation [5, 8, 9, 10].
 In summary, when readmission occurs, it disrupts the care pathway for the patients and poses additional challenges to the clinical team and hospital caring for them.
-Reducing ICU readmission rates might not only improve the care pathway, leading to better patient outcomes but also affect the hospital’s bottom line James (2013); Khera and Krumholz (2018).
+Reducing ICU readmission rates might not only improve the care pathway, leading to better patient outcomes, but also affect the hospital’s bottom line [11, 12].
 Predicting patients at high risk of readmission would not only allow early intervention to reduce the risk but also reduce mortality rates, reduce resource utilisation (based on LOS) and, potentially, hospital costs.
-The emergence of abundant hospital electronic health record (EHR) data Yadav et al.
-(2018); Kim et al.
-(2019); Shickel et al.
-(2017) allowed the use of machine learning (ML) models targeting ICU readmission Badawi and Breslow (2012); Rojas et al.
-(2018); Mohanty et al.
-(2022); Thoral et al.
-(2021); McWilliams et al.
-(2019); Hammer et al.
-(2021); Lin et al.
-(2019); Xue et al.
-(2019); Barbieri et al.
-(2020); Hegselmann et al.
-(2022); Ashfaq et al.
-(2019).
-These readmission models have been summarised in Table S1.
-Apart from their undeniable importance, these models lack proper interpretation, generalisation and validation in different ICU settings (e.g., monocentric versus multicentric environments).
-Our main contention is that predicting ICU readmission in a generalisable way with hospital-based datasets is still a challenge to overcome, even after taking into account these univariate predictive models and best practice ML techniques Lin et al.
-(2019); Hosein et al.
-(2013); Curto et al.
-(2016).
+The emergence of abundant hospital electronic health record (EHR) data [13, 14, 15] allowed the use of machine learning (ML) models targeting ICU readmission [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
+These models for general readmission prediction have been summarised in Table S1.
+Apart from their undeniable importance, these models lack proper interpretation, generalisation and validation in different ICU settings – e.g., monocentric (one hospital centre) versus multicentric (multiple ICU hospital centres).
+Our main contention is that predicting ICU readmission in a well-generali-sable way with hospital-based datasets is still a challenge to overcome, even after taking into account these univariate predictive models and best practice ML techniques [22, 28, 29].
 The main reason for this difficulty is that medical data is usually noisy, uncertain, and characterised by a high degree of missingness, as it relies on different human inputs.
 Further exacerbating this complexity is data imbalance, with only a small proportion of patients being readmitted.
-ICU patients who happen to be readmitted are also very heterogeneous Hosein et al.
-(2013), meaning that finding novel patterns across an ICU readmission data cohort tends to be challenging.
+ICU patients who happen to be readmitted are also very heterogeneous [28], meaning that finding novel patterns across an ICU readmission data cohort tends to be challenging.
 As a result, an appropriate translation from ML models to the clinical environment is still limited.
-In addition, readmission’s association with healthcare quality has been debated in recent years Brown et al.
-(2014); Rewa et al.
-(2018).
-Many readmissions also emerge from factors inherent to patients, not only the quality of care, and may be reduced but not entirely eliminated.
-In these cases, better healthcare quality might not prevent readmission Pakbin et al.
-(2018).
-Still, predicting readmission accurately is valuable for post-discharge monitoring and planning.
-Machine learning models can connect clinical data to readmission, aiding pre-discharge decisions and informing readmission reduction programs Rojas et al.
-(2018); Thoral et al.
-(2021); Huang et al.
-(2021); Hu et al.
-(2022); McIlvennan et al.
-(2015).
-This work proposes a standardised method for a comprehensive assessment of readmission modelling and prediction on ICU data Pollard et al.
-(2018, 2019) with explainable machine learning Molnar (2023).
-Our method resulted in a generalisable predictive model for both multicentric and monocentric data settings Pollard et al.
-(2018, 2019); Johnson et al.
-(2021).
-This is the first study to provide such a large and robust validation on readmission models.
-In addition, explanations from the constructed model also provided a better understanding of the main relevant patient variables associated with 30-day ICU readmission, which can potentially help in guiding clinicians while treating and discharging patients from the ICU.
-1.1 State of Significance
-Problem Intensive Care Unit (ICU) Readmission Prediction with Explainable Machine Learning.
-What is Already Known Although there has been a range of related works aiming to predict ICU readmission with a variable level of success, those studies do not properly assess, characterise and understand readmission prediction, limiting their overall applicability to patient data.
-What This Paper Adds This study proposes a standardised and explainable machine learning (ML) pipeline to model, predict and understand 30-day readmission.
-With a robust validation procedure, the developed ML model achieved generalisable performance on completely new patient data while deriving insightful (literature-supported) conclusions to contrast readmission and non-readmission within an ICU.
+To improve the clinical translation, databases (e.g., MIMIC-IV and eICU) should be homogenised and have their noise and missingness diminished.
+Accordingly, this work proposes a standardised methodological pipeline for a comprehensive assessment of ICU readmission modelling, prediction and understanding [30, 31] using machine learning models [32, 33] and statistical analyses.
+The produced predictive ICU readmission model – which was trained and internally validated in a dataset (i.e., eICU) containing several hospitals (multicentric) and externally validated in a monocentric dataset (i.e., MIMIC IV) [30, 31, 34] – shows clear signs of predictive generalisation.
+It is the first time in the literature that such an approach is employed to confirm the performance of an ICU readmission model.
+To the best of our knowledge, no other previous method has been validated in opposing ICU settings (i.e., monocentric and multicentric).
 2 Material and Methods
-Figure 1 presents the main methodological workflow followed in this work.
-Collecting data from monocentric (i.e., MIMIC IV) and multicentric databases (i.e., eICU) presents itself as the first step.
-Next, from both databases, the creation of variables describing the patient’s information takes place considering their admission timeframe.
-These variables include patient demographics, pre-admission information, vital signs and blood test results, and take into account the first and the last 24 hours, as well as their average values.
-Data preprocessing includes several steps, such as imputation, imbalance learning, and feature standardisation.
-The inclusion of variables outlining the hospitals and ICUs also plays an important role in this characterisation.
-Machine learning algorithms take these features as input to build a readmission model based on a timeframe of 30 days.
-The employed machine learning algorithms use multicentric data from eICU to build and internally validate the predictive model for readmission while it externally tests and validates the final model on monocentric data from MIMIC-IV.
+Figure 1 presents the followed methodological workflow.
+The first step is to collect data from two databases: MIMIC IV (Medical Information Mart for Intensive Care, Version IV) [34, 35] and eICU (eICU Collaborative Research Database) [30, 31, 35].
+From these databases, variables describing the patient’s information are created considering their admission timeframe (see Figure S1).
+These variables include patient demographics, pre-admission information, vital signs and blood test results, and also variables outlining the hospitals and ICUs.
+These variables are fully described in Tables S2 and S3.
+Data preprocessing considers several steps, e.g., imputation, imbalance learning, and feature standardisation.
+Exclusion criteria are also taken into consideration to avoid abnormal patient information in our dataset (see Figure S2).
+Machine learning algorithms take these preprocessed features as input to build a readmission model based on a timeframe of 30 days.
+Hyperparameter optimisation also occurs (Table S4), guaranteeing the selection of the most appropriate predictive model for the data at hand.
+The final employed machine learning algorithms use multicentric data from eICU to build and internally validate the predictive model for readmission while externally testing and validating the final model on monocentric data from MIMIC-IV.
 The readmission model predicts whether a given patient will be readmitted or not and supports explanations for its predictions, driving insights to clinicians while discharging a given patient in the ICU.
 Figure 1: The methodological workflow followed by this work.
-Data comes from monocentric and multicentric databases (i.e., MIMIC IV and eICU, respectively), where the characterisation of ICU patients, hospitals, and ICU takes place.
-Preprocessing filters out, standardises and imputes data for the development of a 30-day readmission machine learning model, which is built and validated on eICU, and tested on MIMIC-IV.
-This model drives explanations from variables and predictions, potentially assisting and guiding clinicians while treating and discharging new patients from the ICU.
-2.1 Data Sources
-For assessing, analysing and understanding ICU readmission, we utilise two datasets: (i) the eICU Collaborative Research Database (eICU) Pollard et al.
-(2018, 2019); Goldberger et al.
-(2000), and (ii) the Medical Information Mart for Intensive Care (MIMIC IV) Johnson et al.
-(2021); Goldberger et al.
-(2000), which are described in next subsections.
-Their access relied on PhysioNet Goldberger et al.
-(2000).
-2.1.1 eICU
-The eICU contains data collected from 335 ICUs at 208 US hospitals in 2014 and 2015.
-This cohort has information about 166,355 unique patients and 200,859 admissions, accounting also for readmissions for those listed patients.
-The eICU cohort is a de-identified multicentric dataset with information about patient demographics, admission and discharge details, APACHE IV scores and variables used for its calculation, vital signs, laboratory test results, medications, and care plan information.
-2.1.2 MIMIC IV
-The MIMIC cohort (version IV) emerged from two in-hospital systems of the Beth Israel Deaconess Medical Center (BIDMC): Electronic Health Record (EHR) and ICU-specific Clinical Information System.
-MIMIC IV has data about patients’ admissions through the emergency department or one of the intensive care units at BIDMC between 2008 and 2019.
-It includes data about 382,278 patients and 523,740 admission records.
-Like eICU, this database is also de-identified and comprises demographics, hospital billing information, data related to patients’ admissions, transfers, and discharges, information about medication administration, and laboratory measurements.
-MIMIC IV also contains data generated in the ICU, such as intravenous and fluid inputs, procedures and services, charted information, and patients’ outputs.
-2.2 Readmission Outcome
-Patients were considered to have met the primary readmission outcome if they had one or more admission episodes within 30 days of discharge from their first ICU admission (see Supplementary Material and Methods for more details).
-If a patient is readmitted twice or more times, only the first readmission within 30 days counts for their data.
-It is worth noting that this range of days to define readmission follows different methodologies Mohanty et al.
-(2022); Lin et al.
-(2019); Xue et al.
-(2019); Barbieri et al.
-(2020).
-Within the 30 days, the eICU cohort had a total of 149,009 patients, where 6,021 and 142,988 were admitted and non-readmitted, respectively.
-On the other hand, MIMIC IV had a total of 68,505 patients, where 5,984 underwent readmission and 62,521 did not.
-2.3 Predictor Variable Selection
-The same potential predictor variables were extracted from both datasets, which resulted in 168 common variables.
-Patient-level data included demographics (age, sex, weight, and height), admission diagnosis, highest and lowest physiological and biochemical values from the first and last 24 hours of ICU admission, treatments provided in ICU (ventilation, dialysis, and vasopressor) and overall illness severity scores (i.e., APACHE IV scores Lee et al.
-(2015)).
-Institutional-level variables included hospital-related (e.g., hospital size and location) and ICU-related variables (e.g., unit type and admission source).
-Supplementary Materials (Table S2) detail and distinguish these common variables in both databases across the following six categories: (i) demographics, (ii) vital sign- and blood test-related, (iii) pathology- and treatment-based, (iv) APACHE severity-related, (v) hospital-related, and (vi) ICU-related.
-In addition, Figure S1 and Supplementary Material and Methods cover more details regarding the data collection criteria and the prediction window.
-Furthermore, categorical variables were encoded in a one-hot encoding scheme (a.k.a, one-of-K or dummy).
-Given the transformation in the categorical feature space, the number of features increased from 168 to 186 variables.
-It is also worth noting that eICU contains an extra level of information on variables when compared to MIMIC IV.
-The variables from the aforementioned categories, which are detailed in Table S2, are only those that match across both cohorts or do not have linkage to the target readmission variable (e.g., length of stay or mortality).
-Nevertheless, as we used eICU to train and validate the machine learning models, we have included important eICU variables (missing in MIMIC IV) to analyse readmission.
-This step has been done only for the sake of readmission explainability.
-This list of complementary variables is detailed in Table S3.
-2.4 Data Preprocessing
-We performed a standardised approach to preprocess both datasets (i.e., eICU and MIMIC IV).
-This approach included basic steps to filter patients’ information and yield valid and reliable information for readmission prediction.
-2.4.1 General Filtering
-First, this step removed patients that are younger than 18 years old from the cohort.
-In addition, the proposed filter did not consider patients who died during their first ICU.
-Furthermore, it separated patients who spent less than four hours in the ICU from the datasets.
-This preprocessing step also excluded patients with abnormal values in the generated variables, such as having an admission weight greater than 260 kg or having an admission height greater than 240 cm.
-Supplementary Material and Methods and Figure S2 detail these patients’ inclusion and exclusion criteria on both eICU and MIMIC IV cohorts.
-2.4.2 Handling Missing Data
-The employed imputation method took into consideration a threshold of 20% of missingness, meaning that our machine learning pipeline will only use variables with less than 20% of missing values.
-The imputation method handled missing values by replacing them with population means (in the case of continuous variables) or a constant (in the case of categorical variables).
-2.4.3 Dealing with Imbalance
-Given the low proportion of readmitted patients in the ICU, the proposed pipeline considered several methods from the Imbalanced-learn toolbox Lemaître et al.
-(2017) to tackle the hardness of learning on imbalanced data Fernández et al.
-(2018).
-Nevertheless, random undersampling over the majority class (i.e., the non-readmitted class) yielded similar results to other methods, being then selected by its simplicity and performance.
-Random undersampling took into accounting for the different data distributions, from 1:1 (1 readmitted patient to 1 non-readmitted patient) to 1:9 (one readmitted patient to 9 non-readmitted patients) to the real distribution (approximately one readmitted patient to 25 non-readmitted patients in eICU, and one readmitted patient to 10 non-readmitted patients in MIMIC IV).
-2.4.4 Standardisation
-In keeping with best practices to improve the predictive performance of variables Raschka et al.
-(2022), standardisation was undertaken by transforming each value of a given variable.
-On each variable value, standardisation subtracts it from the variable’s mean and then divides it by the variable’s standard deviation.
-Although standardisation is not a matter for some machine learning models, it is for several others.
-We therefore applied the same pipeline for all of them, making sure the use of proper data Raschka et al.
-(2022).
-2.5 Machine Learning Modelling, Evaluation and Feature Selection
-Machine learning (ML) modelling followed traditional development and evaluation guidelines Raschka (2018).
-The proposed modelling included the use of seven ensemble-based ML algorithms for patient readmission prediction: (i) explainable boosting machine (EBM) Nori et al.
-(2019), (ii) extremely randomised trees (ERT) Geurts et al.
-(2006), (iii) random forest (RF) Breiman (2001), (iv) extreme gradient boosting (XGBOOST) Chen and Guestrin (2016), (v) adaptive boosting (ADABOOST) Freund and Schapire (1995), (vi) gradient boosting (GB) Friedman (2001), and (vii) light gradient boosting machine (LightGBM) Machado et al.
-(2019).
-Table S4 presents the hyper-parameter optimisations for these ML algorithms.
-It is important to emphasise that these algorithms have been chosen not only because of their robustness based on ensemble properties (i.e., by combining the results of several ML models) but also for their interpretation easiness.
-They all have (at some level) pre-built functions to describe feature importance for a given assessed dataset and can be straightforwardly used with other methods/tools for explainability.
-With these algorithms as base classifiers, a bottom-up greedy feature selection ensued to ensure low ML complexity while maintaining generalisable predictive performance.
-Greedy feature selection starts its iterative process by having a bucket with zero features, adding one feature at each round.
-This method adds a feature in its bucket if it achieves the best predictive performance on a 10-fold cross-validation procedure, which relies on a model built by a classification algorithm.
-The evaluation of each feature relies on an average between Matthew’s correlation coefficient (MCC) Chicco and Jurman (2020) and the area under the receiver operating characteristic (AUC) curve Raschka et al.
-(2022); Raschka (2018); Sokolova and Lapalme (2009), which therefore considers resilience to class imbalance.
-In addition, other evaluation measures complemented the decision process in our ML pipeline, including the area under the precision-recall curve (APR), balanced accuracy, precision, recall (sensitivity), specificity, and F1 score Raschka et al.
-(2022); Raschka (2018); Sokolova and Lapalme (2009).
-2.6 Statistical Analyses
-Data are presented as number/proportion (percentage), mean (± standard deviation) or median (inter-quartile range) depending on the type and distribution of data.
-Chi-squared, ANOVA, Kruskal-Wallis and Wilcoxon rank-sum tests were used to compare groups depending on the type and distribution of data and the number of groups examined Demšar (2006).
-Statistical significance considers p-values lower than 0.001.
-The sample size of eICU and MIMIC IV datasets determined this decision, which consequently guaranteed the robustness of this study.
-It is also worth noting that statistical analyses performed readmission comparisons not only within the same population (e.g., eICU) but also between populations (i.e., eICU versus MIMIC IV), strengthening this methodology.
-Furthermore, APACHE scores are provided only for eICU databases, while not directly available for MIMIC IV, enriching the explanation of characteristics of the former to the latter.
-2.7 Model’s Diagnostic Interpretation, Analysis, and Explanations
-The proposed ML pipeline for readmission also assesses its final model with a posteriori interpretation and explanation methods, which are described next.
-2.7.1 Likelihood Ratios
-Likelihood ratio (LR) McGee (2002) provides a measure for expressing the diagnostic accuracy of a clinical finding of the readmission model from this work.
-LR represents the probability of patients with a condition or undergoing a medical event (e.g., being readmitted) divided by the probability of patients without a condition or undergoing a medical event (e.g., not being readmitted).
-Positive LRs in readmission prediction estimate the model’s capabilities of identifying readmitted against non-readmitted patients.
-When LRs are greater than one, it means they go along with the readmission diagnosis.
-In other words, the greater the LR is, the more certain the predictive model is in finding readmitted patients against non-readmitted patients.
-2.7.2 Calibration
-The calibration technique visually evaluates the predictive readmission model by inspecting the calibration curves and also numerically estimating the calibration slope and intercept, as well as the integrated calibration index (ICI) and associated metrics (E50, E90, Emax ) Austin and Steyerberg (2019).
-The ideal calibration slope is equal to one.
-The ideal calibration intercept, as well as the ideal value for the ICI and its related metrics, is zero.
-2.7.3 Shapley Additive Explanations
-Our pipeline also utilised Shapley Additive Explanations (SHAP) Lundberg and Lee (2017), which is a game theory-based method to derive explanations for individual predictions Molnar (2023), in which combined results yield a global interpretation of the model’s behaviour.
-In readmission prediction, SHAP assists with model analysis, leading to a better understanding of readmission through a range of explanations.
-As a result, it can be used to recommend a set of personalised ICU guidelines while treating/discharging patients in the ICU.
+Data comes from monocentric and multicentric databases (i.e., MIMIC IV and eICU, respectively), where the characterisation of ICU patients, hospitals, and ICU(s) takes place.
+Preprocessing filters out, standardises and imputes data for the development of a 30-day readmission machine learning model, which is built and validated on eICU and tested on MIMIC-IV.
+This model is interpreted and also used to drive explanations from variables and predictions, potentially assisting and guiding clinicians while treating and discharging new patients from the ICU.
+The Supplementary Material (more specifically, Supplementary Methods) provides complete details about data sources (eICU, MIMIC IV, and readmission outcome), predictor variables, data preprocessing, machine learning modelling and validation, and the model’s explanation, diagnosis expression, and calibration.
 3 Results
@@
 Mortality was also slightly higher in the MIMIC IV readmission patients, although no statistical difference has been found between the two databases.
-Among survivors, the readmission group in the eICU was discharged more to home, while readmitted patients in MIMIC IV had a higher rate of further treatments in a rehabilitation care facility (p-value <<< 1%).
+Among survivors, the readmission group in eICU was discharged more to home, while readmitted patients in MIMIC IV had a higher rate of further treatments in a rehabilitation care facility (p-value <<< 1%).
 Table 2: Comparisons of characteristics between eICU and MIMIC IV readmission groups.
@@
 The statistical analysis provided an insightful characterisation of the readmission group, contrasting it with the non-readmission group and across populations.
-With this information, we now move to a machine learning analysis and validation.
-Table S5 shows the 47 selected features by the greedy feature selection approach, considering a balanced readmission dataset.
+With this information, we can now move to a machine learning analysis and validation.
+Table S5 shows the 47 selected features by the greedy feature selection approach using a Random Forest classifier and considering a balanced readmission dataset.
 The details of these variables representing machine learning features can be found in Table S2.
 Although we employed several different undersampling techniques to tackle the high level of readmission imbalance, using a balanced readmission rate presented itself as the most successful technique, which was then applied in both eICU and MIMIC IV cohorts at first.
-Hyper-parameter optimisation has also been applied after feature selection, aiming to improve the predictive performances of the resultant models (Table S4).
+Hyper-parameter optimisation was also applied after feature selection, with the aim of improving the predictive performances of the resultant models (Table S4).
 When bringing all these ML components together, our proposed pipeline achieved an area under the ROC curve (AUC) of 0.68 on multicentric eICU data under a 10-fold cross-validation procedure with a Random Forest classifier with 80 trees in a balanced readmission scenario.
-On the blind test (stratified 10% from eICU data), which internally validates the proposed model, consistent results were reached (AUC of 0.672) when compared to 10-fold cross-validation.
+It is worth noting that this classification model was considered the best against 9680 other machine learning models in terms of a trade-off among AUC, MCC and recall metrics on 10-fold cross-validation.
+Explainability was also a key point in selecting the model.
+Therefore, the selected variables (features) were taken into consideration while selecting the predictive model itself.
+On the blind test (stratified 10% from eICU data), which internally validates the proposed model, consistent results were reached by the predictive model (AUC of 0.672) when compared to 10-fold cross-validation.
 When externally validated on MIMIC IV’s data, our model reached an AUC of 0.616, which demonstrates the overall generalisation capabilities of the proposed model to predict 30-day readmission.
@@
 Figure 2: The predictive performance of the proposed readmission model on 10-fold cross-validation and blind test on eICU data. External validation was made utilising MIMIC IV data.
-3.3 Calibration Analysis
-Calibration metrics are presented in Table S7 for our readmission model on 10-fold cross-validation and blind testing on eICU, and external validation on MIMIC IV.
-Given these metrics, we present the calibration curves for these evaluation sets in Figures S4-S6.
-Figure S4 depicts the calibration curve for our proposed readmission model on 10-fold cross-validation on eICU data.
-The calibration curve on the results in cross-validation indicates a slight overestimation of risk at predicted probabilities between 0.2 and 0.4, and a slight underestimation of risk at predicted probabilities between 0.6 and 0.7.
-Figure S5, in turn, shows the calibration curve for our model on a blind test set from eICU data.
-The calibration curve indicates a slight overestimation of risk at predicted probabilities around 0.5.
-Finally, Figure S6 outlines the calibration curve for our model on an external validation set, i.e., MIMIC IV.
-Its calibration curve indicates a slight underestimation of risk at predicted probabilities below 0.2, and an overestimation of risk at predicted probabilities above 0.4 but most significantly above 0.6.
-Our main conclusion is that our model is well-calibrated in all sets, demonstrating its clinical usability in an ICU setting.
+3.3 Calibration and Likelihood Analysis
+Based on the performed Calibration Analysis (Supplementary Results, Table S7 and Figures S4-S6), the calibration curves, defined from our 30-day ICU readmission model during cross-validation, blind testing and external testing, indicated slight overestimation/underestimation.
+Therefore, our main conclusion is that our model is well-calibrated in all sets, demonstrating its clinical usability in an ICU setting.
 3.4 Likelihood Ratio Analysis
-Figure S7 summarises the likelihood ratios (LRs) for our readmission model, which performs using a classification probability threshold of 0.5.
-With this threshold, the model’s LRs for cross-validation on eICU, blind testing on eICU, and external validation on MIMIC IV are 1.660 (estimated change of probability of 9.628%), 1.636 (estimated change of probability of 9.348%) and 1.335 (estimated change of probability of 5.494%), respectively.
-These model’s results indicate its reasonable diagnostic expression.
-Based on Figure S7, a threshold between 0.70 and 0.82 benefits the model the most regarding the LRs.
-The local and global optima based on the classification threshold 10-fold cross-validation would be 0.712 and 0.813, respectively.
-With 0.712, the model’s LR for 10-fold cross-validation on eICU, blind testing on eICU and external validation on MIMIC IV would be 2.905 (estimated change of probability of 20.26%), 3.326 (estimated change of probability of 22.831%) and 1.942 (estimated change of probability of 12.615%), respectively.
-These estimated changes show small but relevant shifts in the post-test probabilities.
-With 0.813, the model reaches a maximum LR on 10-fold cross-validation on eICU (i.e., 4.256), resulting in a change of probability of readmission for 27.517%.
-Nevertheless, with this threshold, we have a lower LR for blind testing and external validation.
-In the blind test on eICU, an LR of 1.075 would be identified (with an estimated change of probability of 1.373%), while for the external validation on MIMIC IV, the model would result in an LR of 1.400 (with an estimated change of probability of readmission of 6.393%).
-This analysis determines that setting properly the classification threshold is beneficial for having a better diagnostic readmission profile for ICU patients.
-In summary, the model’s diagnostic power depends on a proper analysis comparing and contrasting machine learning and clinical predictiveness.
+In addition, our 30-day ICU readmission model faced an analysis based on Likelihood Ratios (see Supplementary Results and Figure S7), which highlighted the importance of properly setting the classification probability threshold.
+This analysis step is beneficial for having a better diagnostic readmission profile for ICU patients.
+In summary, the ICU predictive model’s diagnostic power depends on a proper analysis of these thresholds, which was properly defined after the Likelihood Ratio Analysis (Supplementary Results).
 3.5 Explainable Machine Learning for ICU Readmission
@@
 Blood Urea Nitrogen (Last 24hs) result are attached to non-readmission.
-In addition, the SHAP tree explanation model resulted in only one hospital-related variable (see Figure 3 and Table S2), i.e., the number of minutes from unit admission time that the patient has been admitted to the hospital (i.e., the Hospital Admission Offset).
-In this case, there is a high correlation between lower values of Hospital Admission Offset and non-readmission, whereas relatively higher values for this variable are more related to readmission.
-From ICU-related variables, the SHAP model outputs a category of unit type as important – i.e., the Unit Type Med-SurgICU, which stands for a unit type of Medical-Surgical ICU.
-As the variable Unit Type corresponds to a categorical variable and Med-SurgICU is one of its options, we transformed its value via one-hot encoding.
-In the SHAP summary plot of Figure 3, this basically means that most patients coming from Med-SurgICU are not facing readmission, while patients that are not passing through such unit type undergo readmission.
-Three important demographics-related variables were also tagged in the SHAP summary from Figure 3: Age, Admission Weight and Admission Height.
-As seen in this figure, older people were often associated with readmission, while low values in age are noticed in non-readmitted patients.
-This factor is actually expected as elderly patients tend to present more comorbidities Divo et al.
-(2014), resulting in more readmissions.
-Contrarily, admission weight followed a distinct pattern, where readmitted patients had low-medium values in their readmission weight.
-We also noticed that readmission and non-readmission were found to be more present in high and low values of the variable admission height, respectively.
-Finally, pathology-related variables, such as the Min.
-Partial Thromboplastin Time (First 24hs), are also at the top 20 features ranked by SHAP values.
-As seen in Figure 3, high Min.
-Partial Thromboplastin Time (First 24hs) values are more connected with readmission.
-We believe these explanations, derived by SHAP for the proposed readmission model, regarding the top 20 most impactful variables might serve as a new source of guidelines while assessing patients through ICU, especially while discharging patients.
-Although clinical practice and experience will always be compulsory in such conditions, we believe machine learning models might provide proper support for them.
 4 Discussion
-4.1 Key Findings
 Although it is challenging to map and match common variables or characteristics across different cohorts, our study drove important clinical insights while analysing eICU and comparing it to MIMIC IV (Section 3.1).
 We discovered how heterogeneous the readmission monocentric and multicentric populations are in general.
-For example, the readmitted patients in MIMIC IV were treated more frequently with ventilation and vasopressors, although renal replacement therapies were more found in patients readmitted to eICU.
+For example, the readmitted patients in MIMIC IV were treated more frequently with ventilation and vasopressors, although renal replacement therapies were more common in patients readmitted to eICU.
 Such contrasts in the cohort populations highlight the ability of our standardised pipeline to deal with heterogeneous readmission data.
@@
 Our proposed machine learning model captured the essence of readmission, being able to transfer similar predictive performance to external validation on monocentric ICU data (Section 3.2).
-Overall, this shows learning on data coming from multiple ICUs has its limitations but may lead to generalisable predictive performance if done adequately.
-A reasonable calibration level is also an essential aspect of the proposed readmission model (Section 3.3).
-As a result, our proposed model shares meaningful clinical decision-making due to its ability to calculate individualised probabilistic estimates of readmission Munn et al.
-(2022).
+Overall, this shows that learning from data coming from multiple ICUs has its limitations but may lead to generalisable predictive performance if done adequately.
+A reasonable calibration level is also an essential aspect of the proposed readmission model (Supplementary Results).
+As a result, our proposed model shares meaningful clinical decision-making due to its ability to calculate individualised probabilistic estimates of readmission [36].
 For example, clinicians may decide that patients with a risk of readmission at least double that of the average risk in their ICU population should be flagged for routine monitoring following discharge from ICU.
@@
 This could potentially facilitate the escalation of ward-based measures to optimise the patient’s management and avoid ICU readmission, or it could facilitate better preparation and planning such that unplanned readmissions are minimised.
-A good diagnostic level, when we increase the classification probability threshold, complements the model’s predictive performance (Section 3.4).
+A good diagnostic level, when we increase the classification probability threshold, complements the model’s predictive performance (Supplementary Results).
 At the threshold(s) of 0.712 (and 0.813), our model returns the most reliable likelihood ratio among cross-validation, blind testing and external validation, yielding trustworthy readmission predictions.
-Finally, we highlight how the 20 most impactful features are linked with readmission prediction through their explainability (Section 3.5).
-These explanations might be used to support the clinical team.
-In fact, some of the insights provided by the interpretation of the readmission model’s explanation are corroborated by the literature.
-We have highlighted these analyses and observations for some of the important variables displayed by SHAP values (Figure 3) next:
-  • 1.
-    Low serum albumin levels in ICU patients have already been linked with death, length of stay, and readmission Rojas et al.
-(2018); Herrmann et al.
-(1992).
-These results from previous work are congruent with our analysis based on Figure 3.
-  • 2.
-    High values of blood urea nitrogen (BUN) have also been associated with readmission and mortality in past studies Rojas et al.
-(2018); Arihan et al.
-(2018); Li et al.
-(2021), representing an important feature to characterise critically ill patients admitted to ICU.
-  • 3.
-    Previous studies also mapped low hemoglobin levels preceding ICU discharge, which may provide important insights for avoiding readmission Hammer et al.
-(2021); Warner et al.
-(2022); Kerfeld et al.
-(2020).
-  • 4.
-    Elderly patients usually are observed to be more associated with ICU readmission Lee et al. (2022); Glans et al. (2020); Pedersen et al. (2017).
-  • 5.
-    The patients’ weight is strongly correlated to readmission.
-While gaining weight through an ICU readmission is linked to non-readmission in patients with heart failure Howie-Esquivel et al.
-(2019), weight and weight change are common factors associated with mortality, duration of mechanical ventilation, and length of stay in the ICU You et al.
-(2013); Mishra et al.
-(2021).
-Therefore, looking at the patient’s admission weight and tracking their respective weight during ICU admission is a relevant aspect before discharging a patient from the ICU.
+Finally, we highlight how the 20 most impactful features are linked with readmission prediction through their explainability (Section 3.3 and Supplementary Results) [17, 37, 38, 39, 21, 40, 41, 42, 43, 44, 45, 46, 47].
 Overall, this shows our proposed model was able to learn and capture good insights on eICU data, being compatible with results and discussion in the ICU readmission literature.
-4.2 Strengths and Limitations
-One of the limitations of this work is the lack of comparable variables between multicentric eICU and monocentric MIMIC IV, including complete APACHE IV scores, which are not present in MIMIC IV.
-Nevertheless, we matched 168 common variables between the two cohorts, constituting both patient- and institutional-level information, building a key strength of our characterisation.
-Another aspect is that several variables presented a large portion of missing data (more than 80%), bringing an extra level of difficulty to perform our methodology.
-We handled missing data by excluding columns with more than 20% of missing values and performing imputation in the other cases.
-This decision was taken because we considered we would include a high proportion of bias and noise if we impute columns with a great frequency of missingness.
-Although missingness might be informative Tan et al.
-(2023), the generalisation reached by the proposed predictive readmission model showed the robustness of our ensemble-based methodology to deal with different noise levels – from both missingness and utilised imputation approach.
-We plan in future work to better understand missingness across eICU and MIMIC IV as a way to improve machine learning modelling.
-In addition, we observed the challenge of finding comparable metrics for the severity of the illness to describe the population between eICU and MIMIC.
-Therefore, our attempt resumed using the variables included in most severity of illness scores in our data description.
-Given data restrictions, we believe this was the best way to compare the severity of illness between both cohorts.
-Finally, we noticed an overall heterogeneity of the studied populations, including imbalance due to rare occurrences of readmission, where the eICU cohort presents 4.04% of readmission cases while MIMIC IV has more than double (8.74%).
-We successfully dealt with the imbalance by employing appropriate performance metrics on top of the feature selection method Chicco and Jurman (2020).
-Our choice for the set of metrics considered the four confusion matrix categories (true positives, false negatives, true negatives, and false positives), setting proportional weights based on the size of the readmission group and the size of the non-readmission group in the datasets.
-Undersampling techniques also assisted with diminishing the imbalance hardness in our study.
-4.3 Relationship to Previous Studies
-Multiple predictive models were developed for readmission with different AUC performances (see Table S1).
-Nevertheless, we believe that most of these models lack clinical applicability.
-The main reason behind this is that those related works prefer highlighting their good predictive performances rather than understanding readmission well and bridging meaningful clinical and machine learning knowledge.
-Although performing similarly to the literature, most of the works are focused on monocentric data, except for Badawi and Breslow Badawi and Breslow (2012) and Mohanty et al.
-Mohanty et al.
-(2022).
-However, both works lack clinical decision-making and are not well-validated with internal and external data sets.
-Apart from Mohanty et al.
-employing explainable machine learning in their proposed models, providing insights for their results, they remain local to the Epic EHR database employed by them.
-Hegselmann et al.’s work Hegselmann et al.
-(2022) is the closest to our method.
-Their readmission model is built from University Hospital Münster (UKM) data and externally tested on MIMIC IV.
-Even though it presents a certain level of similarity to our work, their work is set in a controlled scenario, where UKM data is employed to build the model.
-Furthermore, both UKM and MIMIC IV are monocentric, presenting, therefore, a lower level of difficulty than multicentric eICU.
-In our method, we learn a readmission model from multicentric data to generalise on monocentric data.
-We show our method works well in translating from one to another.
-Funding
+Supplementary Discussion provides more details on the model’s explainability, strength and limitations of this work and its relationships to other studies.
+Authors’ contributions
+Alex G.
+C.
+de Sá: Conceptualisation; Formal analysis; Investigation; Methodology; Project administration; Resources; Software; Validation; Visualisation; Writing – original draft; Writing – review & editing.
+Daniel Gould: Conceptualisation; Formal analysis; Investigation; Methodology; Project administration; Validation; Visualisation; Writing – original draft; Writing – review & editing.
+Anna Fedyukova: Data curation; Formal analysis; Methodology; Software; Validation Visualisation; Writing – original draft.
+Mitchell Nicholas: Validation; Writing – review & editing.
+Calvin Fletcher: Validation; Writing – original draft; Writing – review & editing.
+Lucy Dockrell: Validation; Writing – review & editing.
+David Pilcher: Supervision; Validation; Writing – review & editing.
+Daniel Capurro: Conceptualisation; Validation; Writing – review & editing.
+David B. Ascher: Supervision; Validation; Writing – review & editing.
+Khaled El-Khawas: Conceptualisation; Investigation; Methodology; Project administration; Supervision; Validation; Writing – original draft; Writing – review & editing.
+Douglas E. V. Pires: Conceptualisation; Methodology; Project administration; Supervision Validation; Writing – original draft; Writing – review & editing.
+Acknowledgements
 D.B.
@@
 Supported in part by the Victorian Government’s Operational Infrastructure Support Program.
-Conflict of Interest
+Statement on Conflicts of Interest
 None declared.
-Contributions
-Alex G.
-C.
-de Sá: Conceptualisation, Draft figure proposals, Literature review, Methodology, Machine Learning - Experimental planning and execution, and results analysis, Visualisation, Writing - Original draft preparation and supplementary materials, Reviewing, Editing, and Validation (machine learning perspective);
-Daniel Gould: Conceptualisation, Literature review, Visualisation, Investigation, Writing - introduction and results, Analysis, Reviewing, Editing, and Validation (clinical and machine learning perspectives);
-Anna Fedyukova: Data collection, Machine Learning - Experimental planning and execution, and results analysis, Visualisation, Writing- Methodology and results, supplementary materials, and Validation (machine learning perspective);
-Mitchell Nicholas: Literature review, Reviewing, Editing, and Validation (clinical perspective);
-Calvin Fletcher: Writing - introduction, Validation (clinical perspective), Reviewing, and Editing;
-Lucy Dockrell: Writing, Validation (clinical perspective), Reviewing, and Editing;
-David Pilcher: Supervision, Writing, Validation (clinical perspective), Reviewing, and Editing;
-Daniel Capurro: Conceptualisation, Supervision, Writing, Reviewing and Editing;
-David B. Ascher: Supervision, Writing, Reviewing, Editing and Validation (clinical and machine learning perspectives);
-Khaled El-Khawas: Conceptualisation, Supervision, Supervision, Writing- Methodology and Results, and Validation (clinical perspective);
-Douglas E. V. Pires: Conceptualisation, Supervision, Writing, Reviewing, Editing, and Validation (clinical and machine learning perspectives);
+Summary Table
+  • 1.
+    We propose a standardised and explainable machine learning pipeline for 30-day readmission prediction in multicentric eICU and monocentric MIMIC IV cohorts.
+  • 2.
+    Our random forest-based readmission model yields overall good calibration and generalisable predictive performance on internal and external validation sets.
+  • 3.
+    Calibration and likelihood ratio analysis also demonstrated the proper clinical (diagnostic) usability of our proposed model, even with the challenging nature of the studied readmission datasets.
+  • 4.
+    The derived explanations from the proposed readmission model provide invaluable guidelines, which might be helpful for ICU clinicians’ decision-making while discharging patients.
 Supplementary Information
-Supplementary Materials are available at https://tinyurl.com/36f333sc.
+Supplementary Materials are available at https://tinyurl.com/fne6xv2t.